# Adaptive In-conversation Team Building for Language Model Agents

### v2

@@ 1 Introduction
The success of large language model (LLM) agents (Yao et al., 2022; Yang et al., 2023a; Furuta et al., 2024; Yang et al., 2024a; Hong et al., 2024) with its outstanding in-context learning (Dong et al., 2022; Brown et al., 2020; Yang et al., 2023b; Dai et al., 2023; Li et al., 2023b), planning (Sun et al., 2024; Xie et al., 2024; Liu et al., 2023a; Valmeekam et al., 2022; Wei et al., 2022a; Yuan et al., 2023b; Zheng et al., 2024), tool-using (Qin et al., 2023a; b; Schick et al., 2024; Cai et al., 2023; Yuan et al., 2023a; Paranjape et al., 2023; Zhang et al., 2024b; Huang et al., 2023; Ma et al., 2024), and conversation (Fernandes et al., 2023; Wang et al., 2023c; Yang et al., 2024b) capabilities allow us to relate human’s team building and collaboration abilities to the multiple language model agents (multi-agent) system (Wang et al., 2023a; Xi et al., 2023; Wu et al., 2023; Suzgun & Kalai, 2024a; Hong et al., 2023; Zhang et al., 2024b; 2023a; Valmeekam et al., 2023; Wang et al., 2024; Saha et al., 2023; Liang et al., 2023; Du et al., 2023; Chen et al., 2024). Humans have developed abilities that enable us to form teams and effectively solve problems. These abilities are rooted in communication, social cognition, problem-solving and decision-making, social learning and imitation, and shared intentionality (Elimari & Lafargue, 2020; Confer et al., 2010).
The interplay of the above abilities allows people to organize different teams for problems to ensure that tasks are completed successfully, which brings us to a critical question in a multi-agent system:
@@
A straightforward paradigm would be to build a static agent team beforehand based on the task instruction and let them solve the task collaboratively (Chen et al., 2023; Wu et al., 2023). However, this static build method necessitates maintaining a team with all the required expertise for the whole task cycle. As the complexity of the task increases, the total number of team members may grow significantly. Always proceeding with such a large team makes it challenging to manage the team members effectively and efficiently. Furthermore, static teams may lack the adaptability to respond to dynamic changes in task requirements or unforeseen challenges. Imagine a prehistoric human tribe: was everyone involved in every task? The answer is unlikely affirmative. Those responsible for hunting may not participate in medical care and those responsible for cooking may not involve themselves in management. The major task, survival, was ensured by each individual group sticking to their roles and subtasks. In fact, when human organizations handle a complex task, we tend to form multiple teams for each subtask at different stages of the task-solving procedure, which still guarantees a diverse set of expertise is leveraged demanded by the task complexity (Mao et al., 2016).
@@
We evaluate state-of-the-art multi-agent approaches for complex task solving and our adaptive build approach with Captain Agent on six real-world scenarios, including many mathematics problem-solving (Hendrycks et al., 2021b), data analysis (Hu et al., 2024b), programming (Le et al., 2020), scientific problem-solving (Wang et al., 2023b) (Physics and Chemistry), and world-information retrieval (Mialon et al., 2024). Our experimental results demonstrated the outstanding ability of Captain Agent in various scenarios without heavy prompt engineering for each scenario but only the basic instructions. Captain Agent achieves distinguishing results compared to other single and multi-agent methods and frameworks when using the same prompt for each task, with an average of 21.94% improvement on average accuracy. Ablation studies on static and adaptive building paradigms show that the adaptive team outperforms the static team in four of five scenarios (and matches in one scenario), exhibiting the superiority of the adaptive build paradigm across different scenarios. We also demonstrated that handcraft agents and handcraft tools contribute equally to the final results. We further explore the influence of different backbone LLM for both Captain Agent and nested group chat members or only for nested group chat members. We observe that: (1) Captain Agent with a strong backbone can improve the quality of nested group chat in which the members equipped with weak backbone, and (2) a small model with distinguishable instruction following ability can achieve outstanding performance with low cost.
2 Adaptive In-conversation Team Building
The proposed Captain Agent contains two key components: (1) adaptive multi-agent team-building, which involves agent and tool retrieval, selection, and generation, and (2) nested group conversation with a reflection mechanism within the multi-agent system.
2.1 Overview
The overall workflow of Captain Agent is illustrated in Figure 2. Given a task, Captain Agent is prompted to derive a plan before task execution. According to the plan, Captain Agent will repeat the following two steps until it thinks the task is done and output the results: (Step 1) Captain Agent will first identify a subtask instructed by our prompt, list several roles needed for this subtask, and then create a team of agents accordingly by retrieval, selection, and generation. Each of these will be equipped with predefined tools retrieved from the tool library (Section 2.2); (Step 2) this team of agents will attempt to solve the subtask via conversation with the free-form tool using. Once it’s done, a reflector LLM will provide Captain Agent with a reflection report for it to decide whether to adjust the team or subtask instruction or to terminate and output the results (Section 2.3).
@@ 2.2 Adaptive Multi-agent Team Building
After identifying a subtask in Step 1 following a corresponding prompt, Captain Agent will list several roles for the subtask. These roles will then pass into a retrieval, selection, and generation process guided by Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Ram et al., 2023). Created agents will be equipped with a well-designed profile (system message111System message is used to define an agent’s persona and task-specific instructions.) and high-quality tools. We illustrated the whole process in Figure 3.
Agent and tool retrieval. Captain Agent will prompt required roles with detailed descriptions, including required skills and a possible role name. We use "expert" in Captain Agent prompt to make this process natural. We then retrieve top- agents and top- tools according to the sentence embedding similarity between the role’s description and the agent/tool description recorded in the library. We use Sentence Transformer to calculate the embedding for description between the role and library agents/tools and use cosine similarity as the metric to evaluate the similarity between two sentences, as follows:
where and are the numbers of retrieved agents and tools from agent library and tool library , respectively, for -th role . denotes the sentence embedding extracted from a Sentence Transformer. After retrieval, each role will be assigned with agent candidates and valuable tools. We bind agent candidates with the retrieved tools by injecting the tool-using instruction into the corresponding agent’s system message.
Agent selection. We prompt an LLM-based agent selector to select the most suitable agent according to the role’s description given by Captain Agent and the retrieved agents’ description. A JSON template is designed and provided for the agent selector to ensure the format is correct. Specifically, we designed an abstention mechanism for the agent selector, in which the agent selector can output "None" if there is no suitable agent for a role from the top- retrieved candidate list. This can prevent irrelevant or redundant agents from being forced to be selected for the current task. The roles marked with "None" will further go into the generation process described below.
Agent generation. We design an agent generation process for those roles with no linked agents at the previous step. Specifically, we generate the agent’s name and required skills according to the role description given by Captain Agent. These instructions will be combined with general task and coding instructions and group chat instructions as the final system message. We manually design the general task and coding instructions, motivated by Chain-of-thought (CoT) (Wei et al., 2022b) and Reflexion (Shinn et al., 2024). The final system message will also be compressed to a single-sentence description, which is consumed by the nested group conversation (introduced in the next subsection). We then retrieve tools from the tool library according to the description and inject the tool-using instruction into the generated system message.
Team Memory. Once the team has been built, Captain Agent will cache it into its local memory with a team name and each agent’s detail, including name, system message, and the assigned tools. Captain Agent can call the cached team anytime during the conversation with the user proxy. Calling the cached team will not incur any API calls and thus will not introduce extra costs.
2.3 Nested Group Conversation and Reflection
Agents selected and created in the adaptive multi-agent team-building process will join a nested group chat room. They will be prompted to collect information from the user’s task and solve a subtask from Captain Agent by nested conversation. We then prompt a reflector LLM to retrieve and review the conversation history and fill in the conclusion, the reason for the conclusion, possible contradictions, and issues, and flag if the result needs a double check in the pre-designed template.
Nested group conversation. We perform nested group conversations by leveraging the AutoGen (Wu et al., 2023) framework with a newly designed tool-using paradigm. AutoGen will put all agents in a chat room and select the speaker for each turn by a group chat manager LLM according to the conversation history and each agent’s identity. A short description will be generated from the agent’s profile for the group chat manager. Agents’ code and tool calling will be executed and fed back to the conversation immediately. We inject the tool’s description, path-to-python-module, and response case into the related agent’s system message. The agent can then write free-form code by following the tools’ description and path, naturally incorporating the tools into larger programs. Programs written by all agents will be executed by a user proxy agent with a shared code execution environment, and the results will be fed back to the conversation in real time.
Conversation reflection. The agent’s output during the conversation can be inconsistent, including factual errors, hallucinations, and stereotypes. Although other agents have a chance to adjust and rectify this in conversation, they can also get stuck and cause problem-solving failure. Therefore, we propose to detect such in-conversation contradictions and issues by prompting a reflector LLM with a well-designed conversation summarizing prompt template. The reflector will flag the "need double-check" as "Yes" when it detects such inconsistent content and provides a detailed reason. This will trigger Captain Agent to start a verification process by constructing a new nested conversation to double-check the previous results after receiving "Yes" on "need double-check."
2.4 Benefits over Static Build
A static team with a small number of team members may limit the team’s ability coverage. Although building a large number of agents with comprehensive persona or skill sets can address the limitation in ability coverage, it is challenging for LLMs to handle a long context that introduces all the participant members. Unexpectedly long contexts will primarily reduce the quality of the conversation. Meanwhile, agents with redundant functionality will also be involved in the task-solving process. In contrast, Captain Agent can adaptively select and build more optimized agent teams for the current task, reducing the prompting load for LLMs and redundant output from irrelevant agents without sacrificing the diversity in the agent team.
3 Evaluation
3.1 Experimental Setup
@@
| World Information Retrieval | GAIA (Mialon et al., 2023) | 165 |
@@
Scenarios and datasets. For evaluation, we select various real-world scenarios, including mathematics problem-solving, programming, data analysis, world information retrieval, and science problem-solving. Each scenario was chosen for its unique ability to demonstrate specific capabilities and performance metrics of the agent systems. This ensures a holistic assessment of Captain Agent against the baselines across various critical dimensions of computational and cognitive skills. We bind each scenario with a challenging open-source dataset, as shown in Table 1. Due to cost limitations, we sample a subset of MATH according to its original distribution of each question type.
@@
Compared methods and implementation. For mathematics problems, programming, data analysis, and scientific scenarios, we investigate the performance of Captain Agent and four different methods, including Vanilla LLM (prompt an LLM once for an answer), AutoAgents (Chen et al., 2023), Meta-prompting (Suzgun & Kalai, 2024a), AgentVerse (Chen et al., 2024), DyLAN (Liu et al., 2023b), and a two-agent system (a system involving an Assistant agent with an Executor agent) realized with AutoGen (Wu et al., 2023). Specifically, we implement AutoAgents with AutoGen as the official implementation is unstable and unsuitable for large-scale experiments. For meta-prompting, we improve the code execution ability of meta-prompting by reproducing it with the AutoGen framework. All these methods are equipped with a gpt-4-0125-preview backbone and use the same task-specific prompt (refer to Appendix E).
@@
For world information retrieval scenarios, we compare Captain Agent with the top-5 baselines (with reference) reported to the GAIA validation leaderboard, which includes AutoGen: GAIA_Orchestrator (a specific three-agent setting organized by an Orchestrator agent designed for GAIA) (GAIA_Orchestrator, 2024), FRIDAY (Wu et al., 2024), Warm-up Act222Warm-up Act has no official implementation., and HuggingFace Agent (Huggingface, 2024). All these baselines have a gpt-4-1106-preview backbone, except the HuggingFace Agent equipped with an LLaMA-3-70B as the backbone.
@@
For Captain Agent, we adopt all-mpnet-base-v2 to calculate the sentence embedding for agent and tool retrieval. A User Proxy Agent will communicate with Captain Agent by providing the feedback of code execution, tool calling (adaptive build), nested conversation reflection results, and a default reply: I’m a proxy, and I can only execute your code and tool or end the conversation. If you think the problem is solved, please reply to me only with ’TERMINATE.’
Agent and tool library. We initialize our agent library based on a small subset of problem instances from each dataset (~20 questions per dataset described in Section 3.4) in Table 1. Specifically, we run Captain Agent on the subset and iteratively update the library by adding the generated agents and keeping our agent library unchanged during the main experiment. Our agent library also supports all hand-crafted agents (of the ConversableAgent class) archived in AutoGen (details in Appendix G). All these agents follow the ConversableAgent interface to converse with each other. Our tool library consists of a suite of callable Python functions intended for freeform coding. The agents can freely import functions from the tool library and write free-form code to integrate the outputs to handle sophisticated tasks (see also Appendix F and H). The library contains three main categories of tools: math, data analysis, and world information retrieval. For each category, we summarize the patterns of the corresponding dataset and manually craft a set of functions that suit the tasks.
3.2 Evaluation Protocol
3.3 Main Results
@@
| AutoGPT4 | 13.21 | 0.00 | 3.85 | 4.85 |
| GPT4 Turbo | 20.75 | 5.81 | 0.00 | 9.70 |
| GPT4 + manually selected plugins | 30.30 | 9.70 | 0.00 | 14.6 |
| Captain Agent (Llama-3-70B-Instruct) | 28.30 | 11.63 | 0.00 | 15.15 |
| Huggingface-Agent (Llama-3-70B-Instruct) | 30.19 | 11.63 | 7.69 | 16.97 |
| Warm-up Act | 35.19 | 15.12 | 0 | 17.58 |
| Captain Agent (gpt-4o-mini) | 32.08 | 16.27 | 3.85 | 19.39 |
| FRIDAY | 45.28 | 34.88 | 11.54 | 34.55 |
| AutoGen: GAIA_Orchestrator | 54.72 | 38.31 | 11.54 | 39.39 |
| Captain Agent (gpt-4-0125-preview) | 56.60 | 39.53 | 11.54 | 40.60 |
| Method | Mathematics | Programming | Data Analysis | (Sci) Chemistry | (Sci) Physics |
| Static Team | 64.71 | 88.00 | 85.00 | 47.37 | 68.42 |
| Adaptive Team (Captain Agent) | 82.35 | 96.00 | 95.00 | 52.63 | 68.42 |
| Captain Agent | World-information Retrieval | ||||

### v3

@@ 1 Introduction
The success of large language model (LLM) agents (Yao et al., 2022; Yang et al., 2023a; Furuta et al., 2024; Yang et al., 2024a; Hong et al., 2024) with its outstanding in-context learning (Dong et al., 2022; Brown et al., 2020; Yang et al., 2023b; Dai et al., 2023; Li et al., 2023c), planning (Sun et al., 2024; Xie et al., 2024; Liu et al., 2023a; Valmeekam et al., 2022; Wei et al., 2022a; Yuan et al., 2023b; Zheng et al., 2024), tool-using (Qin et al., 2023a, b; Schick et al., 2024; Cai et al., 2023; Yuan et al., 2023a; Paranjape et al., 2023; Zhang et al., 2024b; Huang et al., 2023; Ma et al., 2024), and conversation (Fernandes et al., 2023; Wang et al., 2023c; Yang et al., 2024b) capabilities allow us to relate human’s team building and collaboration abilities to the multiple language model agents (multi-agent) system (Wang et al., 2023a; Xi et al., 2023; Wu et al., 2023; Suzgun & Kalai, 2024a; Hong et al., 2023; Zhang et al., 2024b, 2023a; Valmeekam et al., 2023; Wang et al., 2024; Saha et al., 2023; Liang et al., 2023; Du et al., 2023; Chen et al., 2024). Humans have developed abilities that enable us to form teams and effectively solve problems. These abilities are rooted in communication, social cognition, problem-solving and decision-making, social learning and imitation, and shared intentionality (Elimari & Lafargue, 2020; Confer et al., 2010). The interplay of the above abilities allows people to organize different teams for problems to ensure that tasks are completed successfully, which brings us to a critical question in a multi-agent system:
@@
A straightforward paradigm would be to build a static agent team beforehand based on the task instruction and let them solve the task collaboratively (Chen et al., 2023; Wu et al., 2023). However, static build necessitates maintaining a team with all the required expertise for the whole task cycle. As the complexity of the task increases, the total number of team members may grow significantly. Always proceeding with such a large team makes it challenging to manage the team members effectively and efficiently. Furthermore, static teams may lack the adaptability to respond to dynamic changes in task requirements or unforeseen challenges. Imagine a prehistoric human tribe: was everyone involved in every task? The answer is unlikely affirmative. Those responsible for hunting may not participate in medical care and those responsible for cooking may not involve themselves in management. The major task, survival, was ensured by each individual group sticking to their roles and subtasks. In fact, when human organizations handle a complex task, we tend to form multiple teams for each subtask at different stages of the task-solving procedure, which still guarantees a diverse set of expertise is leveraged demanded by the task complexity (Mao et al., 2016).
@@
We evaluate state-of-the-art multi-agent approaches for complex task solving and our adaptive build approach with Captain Agent on six real-world scenarios, including many mathematics problem-solving (Hendrycks et al., 2021b), data analysis (Hu et al., 2024b), programming (Le et al., 2020), scientific problem-solving (Wang et al., 2023b) (Physics and Chemistry), and information retrieval (Mialon et al., 2024). Our experimental results demonstrated the outstanding ability of Captain Agent in various scenarios without prompt engineering. Captain Agent achieves distinguishing results compared to other single and multi-agent methods and frameworks when using the same prompt for each task, with an average of 21.94% improvement on average accuracy. Ablation studies on static and adaptive building paradigms show that the adaptive team outperforms the static team in four of five scenarios (and matches in one scenario), exhibiting the superiority of the adaptive build paradigm across different scenarios. We also demonstrated that handcraft agents and handcraft tools contribute equally to the final results. We further explore the influence of different backbone LLM for both Captain Agent and nested group chat members or only for nested group chat members. We observe that: (1) Captain Agent with a strong backbone can improve the quality of nested group chat in which the members equipped with weak backbone, and (2) a small model with distinguishable instruction following ability can achieve outstanding performance with low cost.
2 Captain Agent
The overall workflow of Captain Agent is illustrated in Figure 1. We first establish a main conversation between the user and Captain Agent. When receiving a user task, Captain Agent will: (Step 1) first decompose the task into a list of subtasks and suggest several roles for this subtask. A team of agents will be created accordingly. Each agent will be equipped with several predefined tools according to the abilities defined by Captain Agent (Section 2.1); (Step 2) the agent team will attempt to solve a subtask via a nested conversation. Once the nested conversation reaches the terminal condition, the conversation history will be summarized and analyzed by a reflector LLM, and the final result will be feedback to Captain Agent (Section 2.2). These two steps will repeat certain rounds in the main conversation until reaching the max round limitation or terminal condition.
@@ 2.1 Adaptive Multi-agent Team Building
After identifying a subtask in Step 1 following a corresponding prompt, Captain Agent will list several roles for the subtask. These roles will then pass into a retrieval, selection, and generation process guided by Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Ram et al., 2023). Created agents will be equipped with a well-designed profile (system message111System message refers to a profile that defines an agent’s persona and task-specific instructions.) and high-quality tools.
Agent and tool retrieval. Captain Agent will suggest required roles for the incoming user task at round . includes a short sentence that can describe itself. We then retrieve top- agents and top- tools for according to the embedding similarity between and the agent/tool description in the library, as follows:
where and are the amount of retrieved agents and tools from agent library and tool library , respectively. denotes the extraction of the embeddings. We use to denote the word ”candidate.” We then equip with corresponding .
Agent filtering. Finding an exactly matched agent from our agent library for every real-world task is nearly impossible. Therefore, we prompt an LLM-based filter to select the top- suitable agent for from , simulating the human selection process from massive retrieved results (e.g., searching on the internet). We provide candidate agents , suggested role , and the user task as input to the filter LLM. The filter LLM is required to return the most suitable one agent from . An abstention mechanism is adopted to this process, i.e., the filter LLM can return None to prevent irrelevant agents from being selected as team members. This often happens when is small. We then go through an agent generation process for those with no suitable agents.
Agent generation. We design a step-by-step and templated agent generation process for with no suitable agent selected at the previous step, motivated by Wei et al. (2022b); Shinn et al. (2024). Generally, it includes three steps: (1) generate a basic profile, including name, persona, and required skills as the system message; (2) combine the basic system message with predefined general task-solving and environment interaction instructions, for example, group chat information; (3) retrieved suitable tools from according to the generated system message and bind the new agent with retrieved tools. We then update the newly generated agent into for future use.
Team memory. Captain Agent will cache the built team into its conversation-wise temporary memory, which only includes basic details for each member of the team and no previous conversation history. Captain Agent can prompt the same team with new instructions to seek an alternate result without going through the team-building process, which helps reduce the total cost.
2.2 Nested Conversation and Reflection
Once the team-building process is completed, a new conversation will be established between the created agent team members apart from the main conversation, which we call ”nested conversation.” Captain Agent will provide a subtask to the nested conversation. At the end of the nested conversation, an LLM-based reflector will analyze the whole conversation history. Captain Agent next-step decision (continue or conclude) will based on the reflector’s report.
Nested conversation. In the nested conversation, an LLM-based conversation manager will adaptively select the next speaker from the agent team members based on the conversation history. Team members will share the same conversation history in the nested group chat, but they cannot access the history from the main conversation. Conversation history consists of the previous agent’s action and corresponding feedback. Each agent’s action can be natural language instruction and/or a Python code block. Specifically, we use raw Python function as our tool format, which allows agents to create new tools with their assigned tools by incorporating into a self-written Python code block, which helps extend agent’s action set creatively.
Conversation reflection. A conversation can easily escalate into a conflict once an agent starts a dispute with another, where they will never make a consensus and continue until reaching the max-round. This often happens when Captain Agent provides an ambiguous instruction or the subtask is not clear enough. Therefore, we design a reflection mechanism for nested conversation. An LLM-based reflector will summarize the conversation history and check all disputes inside the conversation, then provide feedback to the Captain Agent with a reflection report. Captain Agent will make further decisions based on this report. All reflection reports will be recorded in the main conversation.
@@ 2.2 Nested Conversation and Reflection
| information retrieval | GAIA (Mialon et al., 2023) | 165 |
@@
2.3 Benefits over Static Team
A static team may limit the team’s ability coverage. Although building a large number of agents with comprehensive persona or skill sets can address the limitation in ability coverage, it is challenging for LLMs to handle a long context that introduces all the participant members. Unexpectedly long contexts will primarily reduce the quality of the conversation. Meanwhile, agents with redundant functionality will also be involved in the task-solving process. In contrast, Captain Agent can adaptively select and build more optimized agent teams for the current task, reducing the prompting load for LLMs and redundant output from irrelevant agents without sacrificing the diversity in the team.
3 Evaluation
3.1 Experimental Setup
@@
Scenarios and datasets. We choose six real-world scenarios, including mathematics, programming, data analysis, information retrieval, and science problem-solving. Each scenario was chosen for its unique ability to demonstrate specific capabilities and performance metrics of the agent systems. This ensures a holistic assessment of Captain Agent against the baselines across various critical dimensions of computational and cognitive skills. We bind each scenario with a challenging open-source dataset, as shown in Table 1. Due to cost limitations, we sample a subset of MATH according to its original distribution of each question type.
@@
Compared methods and implementation. For mathematics problems, programming, data analysis, and scientific scenarios, we investigate the performance of Captain Agent and four different methods, including Vanilla LLM (prompt an LLM once for an answer), AutoAgents (Chen et al., 2023), Meta-prompting (Suzgun & Kalai, 2024a), AgentVerse (Chen et al., 2024), DyLAN (Liu et al., 2023b), and a two-agent system (a system involving an Assistant agent with an Executor agent) realized with AutoGen (Wu et al., 2023). Specifically, we implement AutoAgents with AutoGen as the official implementation is unstable and unsuitable for large-scale experiments. For meta-prompting, we improve its code execution ability by reproducing a real-time Python interpreter. All these methods will be prompted with the same task-specific instruction (refer to Appendix E).
@@
For information retrieval scenarios, we compare Captain Agent with the top-5 baselines (with reference) reported to the GAIA validation leaderboard, which includes AutoGen: GAIA_Orchestrator (a specific three-agent setting organized by an Orchestrator agent designed for GAIA) (GAIA_Orchestrator, 2024), FRIDAY (Wu et al., 2024), Warm-up Act222Warm-up Act has no open-sourced implementation., GAIA official implementation, and AutoGPT (Yang et al., 2023a).
@@
For Captain Agent, we adopt all-mpnet-base-v2 to calculate the sentence embedding for agent and tool retrieval. Our environment will communicate with Captain Agent by providing the feedback of code execution, tool calling (adaptive build), nested conversation reflection results, and a default reply: I’m a proxy, and I can only execute your code and tool or end the conversation. If you think the problem is solved, please reply to me only with ’TERMINATE.’
@@
| AutoGPT (gpt-4) | 13.21 | 0.00 | 3.85 | 4.85 |
| gpt-4-turbo | 20.75 | 5.81 | 0.00 | 9.70 |
| gpt-4-turbo + manually selected plugins | 30.30 | 9.70 | 0.00 | 14.6 |
| Warm-up Act (gpt-4-turbo) | 35.19 | 15.12 | 0 | 17.58 |
| FRIDAY (gpt-4-turbo) | 45.28 | 34.88 | 11.54 | 34.55 |
| AutoGen: GAIA_Orchestrator (gpt-4-turbo) | 54.72 | 38.31 | 11.54 | 39.39 |
| Captain Agent (gpt-4-turbo) | 56.60 | 39.53 | 11.54 | 40.60 |
Agent and tool library. We initialize our agent library based on a small subset of problem instances from each dataset (described in Section 3.4, about 20 questions for each dataset) in Table 1. Specifically, we run Captain Agent on the subset and iteratively update the library by adding the generated agents and keeping our agent library unchanged during the main experiment. Our agent library also supports all hand-crafted agents (of the ConversableAgent class) archived in AutoGen (details in Appendix F). All these agents follow the ConversableAgent interface to converse with each other. Our tool library consists of a suite of callable Python functions intended for freeform coding. The agents can freely import functions from the tool library and write free-form code to integrate the outputs to handle sophisticated tasks (see also Appendix C and G). The library contains three main categories of tools: math, data analysis, and information retrieval. For each category, we summarize the patterns of the corresponding dataset and manually craft a set of functions that suit the tasks.
3.2 Evaluation Protocol
3.3 Main Results
Table 2 and 3 report the comparison results between Captain Agent and eight different baselines on six real-world scenarios. Baseline results on information retrieval are extracted directly from the GAIA leaderboard.
Findings 2: Adaptive team-building boosts performance with no task preference. It is obvious that Captain Agent achieves outstanding results over all scenarios, indicating that Captain Agent is free from task preference. Incorporating different agents into the team at a proper time gives Captain Agent the ability to solve difficult tasks like science and information retrieval problems step-by-step. We observe that Meta-prompting fails in science scenarios due to the inability to decompose science problems into the fine-grain subtasks that one agent can solve. Captain Agent with the agent-team building paradigm neither requires a task that can be decomposed into a subtask that can only be solved by an agent nor requires all agents to be involved in the conversation. Compared with other agent-team building methods such as AutoAgents, DyLAN, and AgentVerse, Captain Agent can not only We further discuss the static and adaptive teams in Section 3.4.1.
3.4 Analysis and Ablation Studies
3.4.1 Static vs. adaptive team-building
To further explore how building a team adaptively in conversation impacts the overall performance, we compare Captain Agent in two settings: (1) building a team at round 1 in the main conversation without changing the team members (static team) and (2) adaptively changing team members and instructions across all rounds in the main conversation (adaptive team). The comparison results in Figure 2 show that except in (Sci) Physics scenario, adaptive build outperforms static build across other scenarios. On data analysis scenarios, adaptive build outperforms static build by over 30.43% on accuracy, showing the importance of changing team members when dealing with data analysis scenarios.
| Captain Agent | Information Retrieval | ||||